OpenStudy (anonymous):

"Please Help" There were 320 students in an auditorium. 3/5 of them were boys. how many more boys than girls were there?

OpenStudy (stacey):

You can find the number of boys by multiplying 320 times 3/5.

OpenStudy (anonymous):

multiply the amount of students by 3/5 in order to find out how many boys the 3/5'ths represent. then once found you can minus that from the original number of students (320). i.e. \[Class = 320\] \[Boys = (3/5)\times320 = 192\] \[Girls = Class - Boys\] \[Girls = 320 - 192 = 108\]

OpenStudy (anonymous):

thank you sssooo much. i really appreciate it.

OpenStudy (stacey):

Girls = 320 - 192 = 128

OpenStudy (anonymous):

ah my bad, what stacey said, i wrote 108 instead of 128 :P

OpenStudy (anonymous):

thank you both sssooo much